caira12:

Solve for x: 1/4 (4x + 15) = 24 (5 points) Select one: a. 49 b. 207/4 c. 81/4 d. 9

Logic007:

Step 1: Multiply both sides by 4

Logic007:

\(4\times\frac{1}{4}(4x+15)=24\times4\)

Logic007:

\(4x+15=96\)

Logic007:

Step 2 : subtract both sides by 15

Logic007:

\(4x+15-15=96-15\)

Logic007:

\(4x=81\)

Logic007:

Step 3 : Divide both sides by 4

Logic007:

\(\frac{4}{4}x=\frac{81}{4}\)

Logic007:

\(x=\frac{81}{4}\)

caira12:

thank you sm

Logic007:

yw! ^